### v1

@@ Abstract
Speculative decoding is an effective and lossless method for Large Language Model (LLM) inference acceleration. It employs a smaller model to generate a draft token sequence, which is then verified by the original base model. In multi-GPU systems, inference latency can be further reduced through tensor parallelism (TP), while the optimal TP size of the draft model is typically smaller than that of the base model, leading to GPU idling during the drafting stage. To solve this problem, we propose EasySpec, a layer-parallel speculation strategy that optimizes the efficiency of multi-GPU utilization. EasySpec breaks the sequential execution order of layers in the drafting model, enabling multi-layer parallelization across devices, albeit with some induced approximation errors. After each drafting-and-verification iteration, the draft model’s key-value (KV) cache is calibrated in a single forward pass, preventing long-term error accumulation at minimal additional latency. We evaluated EasySpec on several mainstream open-source LLMs, using smaller versions of models from the same series as drafters. The results demonstrate that EasySpec can achieve a peak speedup of 4.17x compared to vanilla decoding, while preserving the original distribution of the base LLMs. Specifically, the drafting stage can be accelerated by up to 1.62x with a maximum accuracy drop of only 7%, requiring no training or fine-tuning on the draft models.
@@ 1 Introduction
Figure 1: Overview of EasySpec. (a) Model architecture of layer-parallel speculation. (b) GPU runtime of decoding in the distributed system. SD enables token-level parallelism (fewer gray blocks), while causing multi-GPU under-utilization during the drafting stage. EasySpec solves the problem by layer-parallel speculation.
@@
Transformer-based Large Language Models (LLMs) have demonstrated remarkable problem-solving abilities across various domains Vaswani (2017); Dubey et al. (2024); Yang et al. (2024a; b); Achiam et al. (2023). However, as the parameter size continues to grow, the time-consuming process of auto-regressive decoding poses a significant barrier to deploying large models in latency-sensitive applications Kim et al. (2024); Cai et al. (2024).
@@
Various effective approaches Hinton (2015); Kim et al. (2021); Gu & Dao (2023) have been proposed to reduce inference latency, including speculative decoding Leviathan et al. (2023); Chen et al. (2023) and tensor-parallel (TP) distribution Shoeybi et al. (2019). Speculative decoding employs a smaller model to generate a draft token sequence, and uses token-level parallelism to conduct non-autoregressive verification, ensuring no shifting of the original model’s output distribution. In contrast, TP distribution leverages cross-device parallelism by partitioning computational workloads across multiple devices (usually GPUs) and synchronizing the results subsequently, which is also lossless.
@@
Combining speculative decoding with TP distribution achieves even greater acceleration rates. However, integrating a draft model into a distributed system is not trivial Cai et al. (2024). Since the parameter size of the draft model is typically much smaller (around or less than 1/10) than that of the base model, the optimal TP size (the number of workload segments) is correspondingly smaller Chen et al. (2023), meaning that the draft model would run fastest when dispatched on one or a subset of GPUs, leaving other GPUs idle (see Figure 1 and Table 1). Consequently, multi-GPU computational resources are under-utilized during the drafting stage.
@@
We identify the primary cause of such inefficiency as the lack of parallelism between the draft model’s layers: while tensor operations within one layer can be parallelized by TP, the layers themselves are restricted to be executed sequentially, one after another and from bottom to top, for generating a “precise” result of inference. However, the drafting result is never required to be precise, as it is only used for token parallelism and does not directly impact the final output—the verification result does (see Section 2.1). Therefore, strictly following the execution order is unnecessary, while a “fuzzy” but faster layer execution strategy could be preferable than the precise one, as long as it can sufficiently approximate the drafting result.
@@
Based on this insight, we propose EasySpec, a layer-parallel speculation strategy for optimizing the efficiency of drafting-stage multi-GPU utilization. EasySpec introduces fuzzy speculation by breaking the restriction of the execution order and running multiple attention layers simultaneously, with the most recent hidden state as input to all of them (Figure 1). This layer-level parallelism enables more devices to participate in the drafting stage, as the data dependencies between some consecutive layers are removed (Algorithm 2).
@@
Fuzzy speculation may introduce approximation errors, as the inference procedure differs from how the model is pre-trained. To prevent long-term error accumulation in the key-value (KV) cache of the draft model, we perform a process called bonus calibration after each iteration of drafting-and-verification. Firstly, the accepted tokens are concatenated to the bonus token to form a token sequence. This sequence is then re-input to the draft model, where a standard layer-sequential forward pass is executed, and the KV cache is updated with the precise values. This calibration step applies token parallelism, which is typically used only in the verification stage, to the drafting stage as well. Therefore, it incurs minimal additional latency—roughly equivalent to that of a standard forward pass.
@@
We evaluate our method on two widely-used open-source LLMs, Llama-3-70B-Instruct and Qwen2-72B-Instruct, as well as a state-of-the-art LLM, Llama-3.3-70B-Instruct, and two task-specific models Qwen2-Math-72B-Instruct and Qwen2.5-Coder-32B-Instruct Hui et al. (2024). The draft models are selected from the same series as their corresponding large models. The evaluation results show that our method can achieve a peak speedup of 4.17x over vanilla decoding. Specifically, the drafting stage can be accelerated by up to 1.62x with no greater than 7% drop of accuracy, requiring no additional training or fine-tuning on the existing draft models.
@@ 2.1 Speculative decoding
Through this whole verification process, the final output token sequence aligns with the original distribution of the base model. Details of this algorithm are in Appendix A.
@@ 2.2 Tensor-Parallel Distributed Inference
Tensor parallelism is a technique for deploying large models across multiple devices (GPUs). As for TP distributed inference, tensors are split up into multiple chunks, and each of these chunks is dispatched on a unique GPU. During computation operations, each GPU works on its respective chunk independently at first, and the results are synchronized afterwards to ensure that the final result is identical to the original. In modern multi-GPU systems, advanced communication hardware and protocols enable TP to be applied for inference acceleration.
@@ 2.3 Device Characteristics
Modern GPUs feature powerful computation capacities and hence short computation latency. Therefore, matrix computation involved in LLM inference is usually bottlenecked by memory bandwidth or control-flow overheads Kim et al. (2023). In multi-GPU systems, it can also be bounded by communication and synchronization overheads.
Formally, we denote the workload of 1-token forward as $W$, the latency of executing $W$ on a single GPU as $T_{single}(W)$, and the additional overhead of TP distribution as $T_{add}$. The workload of $s$-token parallelism is $sW$, and $W/s$ the workload of $s$-sized TP. If $W$ is relatively small compared to the computation power, we can expect the following relationships:
$T_{single}(W)\approx T_{single}(sW)<sT_{single}(W)$ (2)
$T_{single}(W/s)+T_{add}>T_{single}(W)$ (3)
The effectiveness of token-level parallelism can be illustrated by Equation 2, which is leveraged by speculative decoding for acceleration. Equation 3 highlights the impact of communication and control-flow overheads in distributed inference, which constrains multi-GPU utilization of the draft model.
@@ 3.1 Fuzzy Speculation
As mentioned in Section 1, the standard speculation process restricts the layers to be executed sequentially. Algorithm 1 demonstrates the execution procedure of $N$ consecutive layers. If we eliminate $h^{\prime}_{i}$ in the algorithm, we have
@@
Algorithm 1 Sequential Precise Speculation
  Input: hidden state $h$, $N$ consecutive attention layers $Attn_{1}$$\cdots$,$Attn_{N}$ and MLP layers $MLP_{1}$$\cdots$,$MLP_{N}$
@@
  Input: hidden state $h$, $N$ consecutive attention layers $Attn_{1}$$\cdots$,$Attn_{N}$ and MLP layers $MLP_{1}$$\cdots$,$MLP_{N}$
@@
We propose an effective approximation method for fuzzy speculation. Inspired by Lee et al. (2024), we observe that the attention layers’ outputs can be well approximated by substituting the input $h_{i}$ with $h_{1}$ for all $1<i\leq N$. Algorithm 2 shows the detailed steps of this method. From the perspective of model architectures, the input of the attention layers is not the most recent hidden states, but rather those from several layers prior (Figure 1). With this modification of layer execution strategy, we break the data dependencies between $N$ consecutive attention layers: once $h_{j}$ is computed for some $j$, all attention layers between $j$ to $j+N$ can be executed with no data dependencies among them, thereby achieving a layer-level concurrency. In a multi-GPU system, these attention layers can run simultaneously on different devices (Figure 1).
@@
$\begin{split}T_{seq}=NT_{single}(A)\\ T_{fuzzy}=T_{single}(A)+T_{add}\end{split}$ (5)
where $A$ is the workload of one attention layer. Unlike the traditional TP method, the latency of running an attention layer is long enough to compensate for the additional overhead, that is $(N-1)T_{single}(A)>T_{add}$ for $N>1$, resulting in $T_{fuzzy}<T_{seq}$. The MLP layers are executed in the original sequential order, while we can run identical copies of them on all devices to eliminate communication, fully utilizing the multi-GPU computational resources.
Fuzzy speculation will inevitably result in some loss of speculation accuracy. In order to maintain a higher acceptance rate, EasySpec incorporates tree attention Miao et al. (2024), a technique commonly used in existing speculative decoding methods Cai et al. (2024); Li et al. (2024a; b). Tree attention can boost the acceptance rate by increasing the number of draft token sequences. As shown in Figure 2, the drafted tokens are structured as a tree, with each path corresponding to a sequence. With 2 forward passes, 6 draft token sequences are generated. Speculation and verification for all the sequences can be performed in a single forward pass using a 2D tree-attention mask.
Figure 2: Illustration of tree attention.
Figure 3: Illustration of bonus calibration.
@@ 3.2 Bonus Calibration
The precise values of $kv^{\prime}$ are required for preventing such error accumulation. Nevertheless, obtaining these values demands running the model’s layers sequentially, which is exactly what we try to optimize through fuzzy speculation. To solve this dilemma, we propose an efficient way of obtaining precise KV values with minimal additional latency. At the beginning of a standard drafting stage, the bonus token $x_{t+m+1}$ from the last iteration is input to the draft model, for the speculation of $x^{\prime}_{t+m+2}$. This is a 1-token forward pass and also memory-bounded, as discussed in Section 2.1. Therefore, similar to the verification stage, token parallelism can also be applied here. By concatenating the accepted token sequence $x_{t+1},\cdots,x_{t+m}$ with the bonus token $x_{t+m+1}$ and re-inputting this entire sequence $x_{t+1},\cdots,x_{t+m},x_{t+m+1}$ to the draft model, the precise KV values $kv^{\prime}_{t+1},\cdots,kv^{\prime}_{t+m}$ (and $kv^{\prime}_{t+m+1}$) can be produced through a single sequential forward pass. According to Equation 2, the latency of such a forward is nearly equal to a 1-token sequential forward.
We propose bonus calibration based on the method above. As shown in Figure 3, we firstly discard all fuzzy KV items, regardless of whether the corresponding tokens are accepted or rejected (e.g. the green items 1 and 2 are also discarded). Then, a token-parallel forward is conducted to refill the KV cache with precise values. From the perspective of execution, the precise KV values are by-products of the next-token speculation of the bonus token, similar to how the bonus token itself is produced by the verification. This is the reason why we call this process “bonus” as well. Bonus calibration can generate a precise candidate token and distribution, while subsequent rounds of speculation can revert to be fuzzy. The overall latency increases slightly, from $nT_{fuzzy}$ to $T_{seq}+(n-1)T_{fuzzy}$ ($n$ is the speculation length), while the elimination of errors is crucial and substantial.

### v2

@@ Abstract
Speculative decoding is an effective and lossless method for Large Language Model (LLM) inference acceleration. It employs a smaller model to generate a draft token sequence, which is then verified by the original base model. In multi-GPU systems, inference latency can be further reduced through tensor parallelism (TP), while the optimal TP size of the draft model is typically smaller than that of the base model, leading to GPU idling during the drafting stage. We observe that such inefficiency stems from the sequential execution of layers, which is seemingly natural but actually unnecessary. Therefore, we propose EasySpec, a layer-parallel speculation strategy that optimizes the efficiency of multi-GPU utilization. EasySpec breaks the inter-layer data dependencies in the draft model, enabling multiple layers to run simultaneously across multiple devices as “fuzzy” speculation. After each drafting-and-verification iteration, the draft model’s key-value cache is calibrated in a single forward pass, preventing long-term fuzzy-error accumulation at minimal additional latency. EasySpec is a training-free and plug-in method. We evaluated EasySpec on several mainstream open-source LLMs, using smaller versions of models from the same series as drafters. The results demonstrate that EasySpec can achieve a peak speedup of 4.17x compared to vanilla decoding, while preserving the original distributions of the base LLMs. Specifically, the drafting stage can be accelerated by up to 1.62x with a maximum speculation accuracy drop of only 7%. The code is available at https://github.com/Yize-Wu/EasySpec.
Figure 1: Overview of EasySpec: comparison between standard Speculative Decoding and EasySpec. The differences are mainly about: (1) Fuzzy Speculation uses layer parallelization for drafting acceleration. (2) The fuzzy KV items are discarded regardless of acceptance length, while Speculative Decoding preserves accepted items. (3) Speculative Decoding is 2-staged and EasySpec is 3-staged, with an additional stage for bonus calibration.
Figure 2: GPU runtime of different decoding algorithms in the distributed system. Speculative Decoding (SD) enables token-level parallelism (fewer base-model blocks), while causing multi-GPU under-utilization during the drafting stage. EasySpec solves the problem by layer-parallel speculation.
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) have demonstrated remarkable problem-solving abilities across various domains [vaswani2017attention, dubey2024llama, yang2024qwen2technicalreport, yang2024qwen2, achiam2023gpt]. However, as the parameter size continues to grow, the time-consuming process of auto-regressive decoding poses a significant barrier to deploying large models in latency-sensitive applications [kim2024speculative, cai2024medusa].
@@
Various effective approaches [hinton2015distilling, kim2021bert, gu2023mamba] have been proposed to reduce inference latency, including speculative decoding [leviathan2023fast, chen2023accelerating] and tensor-parallel (TP) distribution [shoeybi2019megatron]. Speculative decoding employs a smaller model to generate a draft token sequence, and uses token-level parallelism to conduct non-autoregressive verification, ensuring no shifting of the original model’s output distribution. In contrast, TP distribution leverages cross-device parallelism by partitioning computational workloads across multiple devices (usually GPUs) and synchronizing the results subsequently, which is also lossless.
@@
Combining speculative decoding with TP distribution achieves even greater acceleration ratio. However, integrating a draft model into a distributed system is not trivial [cai2024medusa]. Since the parameter size of the draft model is typically smaller than that of the base model, the optimal TP size (the number of segments of workload distribution) is correspondingly smaller [chen2023accelerating], meaning that the draft model would run fastest when dispatched on one or a subset of GPUs, leaving other GPUs idle (see Figure˜2 and Table˜6). Consequently, multi-GPU computational resources are under-utilized during the drafting stage.
@@
We identify the primary cause of such inefficiency as the lack of parallelism between the draft model’s layers: while tensor operations within one layer can be parallelized by TP, the layers themselves are restricted to be executed sequentially, one after another and from bottom to top, for generating a “precise” result of inference. However, the drafting result is never required to be precise, as it is only used for token parallelism and does not directly impact the final output—the verification result does (see Section˜2.1). Therefore, strictly following the execution order is unnecessary, while a “fuzzy” but faster layer execution strategy could be preferable than the precise one, as long as it can sufficiently approximate the drafting result.
@@
Based on this insight, we propose EasySpec, a layer-parallel speculation strategy for optimizing the efficiency of drafting-stage multi-GPU utilization. EasySpec introduces fuzzy speculation, which breaks the data dependencies between some consecutive attention layers by running them with the most recent hidden state as input to all of them (Figure˜1 and Algorithm˜2). As the data dependencies are eliminated, multiple layers of the draft model can execute simultaneously on separate devices, achieving layer-level parallelism, more efficient multi-GPU utilization and hence faster speculation inference. Meanwhile, the speculation outputs can be well-approximated, retaining a high speculation accuracy.
@@
The approximation errors in the key-value (KV) cache of the draft model may accumulate during the inference procedure. To prevent this, we perform bonus calibration after each iteration of drafting-and-verification. Firstly, the accepted tokens are concatenated to the bonus token to form a token sequence. This sequence is then re-input to the draft model, where a conventional layer-sequential forward pass is executed, and the KV cache is updated with the precise values. This calibration step applies token parallelism, which is typically used only in the verification stage, to the drafting stage as well. Therefore, it incurs minimal additional latency to the speculation procedure, while significantly enhancing accuracy—ultimately contributing to improved overall inference speed.
@@
We evaluate our method on several widely-used open-source LLMs and task-specific models. The draft models are selected from the same series as their corresponding large models. The evaluation results show that EasySpec can achieve a peak speedup of 4.17x over vanilla decoding. Specifically, the drafting stage can be accelerated by up to 1.62x with no greater than 7% drop of speculation accuracy, requiring no additional training or fine-tuning on the existing draft models.
@@ 2.1 Speculative decoding
Through this whole verification process, the final output token sequence aligns with the original distribution of the base model. Details of this algorithm are in Appendix˜A.
@@ 2.2 Tensor-parallel distributed inference
Tensor parallelism is a technique for deploying large models across multiple devices (GPUs). As for TP distributed inference, tensors are split up into multiple chunks, and each of these chunks is dispatched on a unique GPU. During computation operations, each GPU works on its respective chunk independently at first, and the results are synchronized afterwards to ensure that the final result is identical to the original. In modern multi-GPU systems, advanced hardware and communication protocols enable TP to be applied for inference acceleration.
The optimal TP size is primarily influenced by the parameter size. Firstly, the TP size must evenly divide both the hidden dimension and the number of attention heads. For example, in Qwen-2-0.5B-Instruct these numbers are 4864 and 14 respectively, which limits the maximum feasible TP size to 2. Secondly, the computational workload varies with parameter size, which further impacts TP efficiency. A detailed discussion of these device-level characteristics is provided in Section˜2.3.
@@ 2.3 Device characteristics
Modern GPUs feature powerful computation capacities and hence short computation latency. Therefore, matrix computation involved in LLM inference is usually bottlenecked by memory bandwidth or control-flow overheads [kim2023squeezellm]. In multi-GPU systems, the computation can also be bounded by communication and synchronization overheads.
Formally, we denote the workload of 1-token forward as $W$, the latency of executing $W$ on a single GPU as $T_{exe}(W)$, and the additional overhead of TP distribution as $T_{addi}$. The workload of $s$-token parallelism is $sW$, and $W/s$ the workload of $s$-sized TP. If $W$ is relatively small compared to the computation power, we can expect the following relationships:
$T_{exe}(W)\approx T_{exe}(sW)<sT_{exe}(W)$ (2)
$T_{exe}(W/s)+T_{addi}>T_{exe}(W)$ (3)
The effectiveness of token-level parallelism can be illustrated by Equation˜2, which is leveraged by speculative decoding for acceleration. Equation˜3 highlights the impact of communication and control-flow overheads in distributed inference, which constrains the TP size not to be too large and limits multi-GPU utilization of the draft model.
@@ 3.1 Fuzzy speculation
As mentioned in Section˜1, the conventional speculation process restricts the layers to be executed sequentially. Algorithm˜1 demonstrates the execution procedure of $N$ consecutive layers. If we eliminate $h^{\prime}_{i}$ in the algorithm, we have
@@
Algorithm 1 Layer-Sequential Speculation
[…]
@@
[…]
@@
Figure 3: 2-D demonstration of high cosine similarity between $h_{i}$ and $h_{i+1}$.
We propose an effective approximation method for fuzzy speculation. Inspired by [lee2024infinigen], we observe that the attention layers’ outputs can be well approximated by substituting the input $h_{i}$ with $h_{1}$ for all $1<i\leq N$. In Equation˜4, the hidden state update $h_{i+1}-h_{i}=Attnoutput_{i}+MLPoutput_{i}$ is relatively less significant, so the cosine similarity between $h_{i}$ and $h_{i+1}$ can approach $1$, and consequently the similarity between all $h_{i}$ and $h_{1}$ is also high (Figure˜3 is a 2-D demonstration of this phenomenon). Such high similarities of the input allow other hidden states to be accurately approximated as well (see Section˜E.1 for theoretical explanations and Table˜4 for empirical results).
Algorithm˜2 shows the detailed steps of this method. From the perspective of model architectures, the input of the attention layers is not the most recent hidden states, but rather those from several layers prior (Figure˜1). With this modification of layer execution strategy, we break the data dependencies between $N$ consecutive attention layers: once $h_{j}$ is computed for some $j$, all attention layers between $j$ to $j+N$ can be executed with no data dependencies among them, thereby achieving a layer-level concurrency. In a multi-GPU system, these attention layers can run simultaneously on different devices (Figure˜2).
@@
$\begin{split}T_{seq}=NT_{exe}(A)\\ T_{fuzzy}=T_{exe}(A)+T_{addi}\end{split}$ (5)
where $A$ is the workload of one attention layer. Unlike the traditional TP method, the latency of running an attention layer is long enough to compensate for the additional overhead, that is $(N-1)T_{exe}(A)>>T_{addi}$ for $N>1$, resulting in $T_{fuzzy}<T_{seq}$. The MLP layers are executed in the original sequential order, as parallelizing them results in significant approximation errors and degraded end-to-end performance (Appendix˜B).
In order to achieve a higher acceptance rate, EasySpec also incorporates tree attention [miao2024specinfer] in the drafting stage, a technique commonly used in existing speculative decoding methods [cai2024medusa, li2024eagle, li2024eagle2]. Tree attention can boost the acceptance rate by increasing the number of draft token sequences. As shown in Figure˜5, the drafted tokens are structured as a tree, with each path corresponding to a sequence. Speculation and verification for all the sequences can be performed in a single forward pass using a 2D tree-attention mask.
3.1.1 Layer-parallel strategy
For the layer-parallel size $N$, we exclude the first and last layer and parallelize $1\sim N-1$, $N\sim 2N-1$, …, until the remaining layers (except for the last) are not enough to compose a $N$-set. For instance, for Llama-3-8B-Instruct (32 layers) and $N=4$, we have 1 $\sim$ 3, 4 $\sim$ 7, …, 28 $\sim$ 30 parallelized, while layer 28,29, and 30 have to compose a 3-set because layer 31 is the last (see Table˜1). The reason of such strategy is that, parallelizing intermediate layers (excluding the first and last layer) yields best performances in our experiments. We hypothesize that this is because the first and last layer are more sensitive to approximation errors.
Note that there could be a better strategy for models and/or tasks, but this strategy has already yielded sufficiently satisfactory acceleration. Moreover, trying other strategies is simple and lightweight, as it requires no modifications to the original model architecture. A change in the configuration is enough.
Table 1: Example layer-parallel strategies for drafter models. The strategies for other drafters is the same as specified in Section˜3.1.1.
Model # of Layers NN Strategy
Llama-3-8B-Instruct 32 4 0, 1-3, 4-7, 8-11, …, 24-27, 28-30, 31
Llama-3.2-1B-Instruct 16 2 0, 1-2, 3-4, 5-6, …, 11-12, 13-14, 15
Qwen2-7B-Instruct 28 4 0, 1-3, 4-7, 8-11, …, 20-23, 24-26, 27
@@ 3.2 Bonus calibration
The precise values of $kv^{\prime}$ are required for preventing such error accumulation. Nevertheless, obtaining these values demands running the model’s layers sequentially, which is exactly what we try to optimize through fuzzy speculation. To solve this dilemma, we propose an efficient way of obtaining precise KV values with minimal additional latency. At the beginning of a standard drafting stage, the bonus token $x_{t+m+1}$ from the last iteration is input to the draft model, for the speculation of $x^{\prime}_{t+m+2}$. This is a 1-token forward pass and also memory-bounded, as discussed in Section˜2.1. Therefore, similar to the verification stage, token parallelism can also be applied here. By concatenating the accepted token sequence $x_{t+1},\cdots,x_{t+m}$ with the bonus token $x_{t+m+1}$ and re-inputting this entire sequence $x_{t+1},\cdots,x_{t+m},x_{t+m+1}$ to the draft model, the precise KV values $kv^{\prime}_{t+1},\cdots,kv^{\prime}_{t+m}$ (and $kv^{\prime}_{t+m+1}$) can be produced through a single sequential forward pass. According to Equation˜2, the latency of such a forward is nearly equal to a 1-token sequential forward.
We propose bonus calibration based on the method above. As shown in Figure˜1, we firstly discard all fuzzy KV items, regardless of whether the corresponding tokens are accepted or rejected (e.g. the accepted items 1 and 2 are also discarded). Then, a token-parallel forward is conducted to refill the KV cache with precise values. From the perspective of execution, the precise KV values are by-products of the next-token speculation of the bonus token, similar to how the bonus token itself is produced by the verification. This is the reason why we call this process “bonus”. Bonus calibration can generate a precise candidate token and distribution, while subsequent rounds of speculation can revert to be fuzzy. The overall latency increases slightly, from $nT_{fuzzy}$ to $T_{seq}+(n-1)T_{fuzzy}$ ($n$ is the speculation length), while the elimination of errors is crucial and substantial. More detailed theoretical backing is in Section˜E.2.